# Cross-Modal Geometric Hierarchy Fusion: An Implicit-Submap Driven Framework for Resilient 3D Place Recognition

### v1

@@ III-B Implicit Submap Representation With Elastic Neural Points
We will employ a memory-efficient implicit neural representation to store the geometry of the sub-map. Based on elastic neural points, this representation possesses local rotational and translational invariance. Additionally, it is unaffected by the density of the original input point cloud, has a uniform distribution, and can remove interference from certain dynamic objects.
  1. 1)
    Elastic Implicit Neural Point Cloud: The results of the sub-map composed of neural implicit points $\mathcal{M}$ are shown as follows:
    =$\mathcal{M}=\left\{\mathbf{m_{i}}=\left(\mathbf{p_{i}},\mathbf{q_{i}},\mathbf{f_{i}},t_{i}^{c},t_{i}^{u},\mu_{i}\right)\mid i=1,\dots,N\right\}$ (1)
@@
  2. 2)
    SDF Decoder: The joint feature coordinate encoding for scene representation is employed to achieve fast convergence and reconstruct surface holes. To ensure invariance of the prediction to the local translation and rotation of neural points, the query position is transformed from the global frame to the local frame of each neural point. Specifically, let $\mathbf{d_{j}}\in\mathbb{R}^{3}$ denote the query position in the local frame of neural point $\mathbf{m_{j}}$. This local coordinate $\mathbf{d_{j}}$ is obtained by transforming the global coordinates $\mathbf{p}_{s}$ of the query position using the pose of neural point $\mathbf{m_{j}}$, as follows:
@@
We employ a dynamic point filtering method based on the stability $\mu_{i}$ and SDF values of neural points. When updating the submap, a point is classified as a dynamic point and filtered out if its SDF value exceeds a dynamic distance threshold $\gamma_{d}$ and its stability value exceeds a stability threshold $\gamma_{\mu}$, Specifically, a sampled point $\mathbf{p_{W}}$ in the world coordinate system is considered dynamic if it satisfies the condition $\mathbf{S}(\mathbf{p_{W}})>\gamma_{d}$ and $\mathbf{H}(\mathbf{p_{W}})>\gamma_{\mu}$, where $\mathbf{S}(\mathbf{p_{W}})$ represents the SDF value indicating the distance from the point to the nearest surface, and $\mathbf{H}(\mathbf{p_{W}})$ represents the stability value quantifying the point’s consistency across multiple observations. This method effectively removes dynamic objects from the map, enhancing its accuracy and consistency.
@@ III-C Normal Vector Descriptor
After obtaining the implicit representation of the submaps, we extract the scene’s occupancy information, the mesh vertices, and their corresponding normal vectors.
The process begins by discretizing the submap into a voxel grid with a predefined resolution $r$. At each voxel corner point within this grid, we query the Signed Distance Function (SDF) value according to the method described previously. Using these SDF values, we apply an enhanced Marching Cubes (MC) algorithm to reconstruct a triangular mesh. The MC algorithm generates vertices uniformly distributed on the reconstructed surface. The normal vector is then calculated for each vertex, defining the local surface orientation. Next, we filter the vertices to remove isolated outliers. This step refines the mesh by retaining only significant vertices along with their normals. Finally, occupancy information is derived by analyzing the SDF values at the corners of each voxel. Voxels exhibiting a sign change in their corner SDF values are classified as occupied. This occupancy classification is essential for distinguishing between occupied and free space in the environment.
We use the GPF [72] algorithm to filter out the road surface points from the input mesh vertices. This is because the road surface points contain little geometric information and provide little benefit for extracting feature descriptors for place recognition. Subsequently, inspired by the work of FEC [73], we apply a fast clustering method for mesh vertices. This method takes the mesh vertices from which the ground points have been removed as input, and the algorithm is as follows:
@@
1:mesh vertices point cloud $\mathbf{cloud}$, minimum cluster size $min\_size$, distance tolerance $tol$
@@
4:$n\leftarrow|\mathbf{cloud}|$
5:if $n<min\_size$ then
@@
15:     if $neighbors$ is empty then
@@
29:         if $labels[j]>min\_tag$ then
30:              Relabel all points with $labels[k]=labels[j]$ to $min\_tag$
@@
Following the derivation of clusters $\mathbf{\mathcal{C}=\left\{C_{1},...,C_{M}\right\}}$, points within each cluster are normalized. This normalization reduces sensitivity to noise and improves the robustness of subsequent geometric feature extraction. Specifically, all points in a cluster are isotropically scaled to be contained within a unit sphere of radius 1.
@@
Next, the spherical domain is partitioned into 72 sub-regions based on longitude and latitude. The sphere is first divided along the z-axis into northern and southern hemispheres. By latitude, each hemisphere is further divided into two bands, resulting in 4 latitude bands globally. Longitudinal division then subdivides each latitude band into 18 sectors, formed by meridians spaced 20 degrees apart. Fig. 3 illustrates this spatial subdivision.
@@
Figure 3: The figure illustrates the extraction of geometric descriptors for 3D segments. The input is a 3D segment with surface normals; color coding indicates normal deviations. First, these normals are mapped to discrete bins on a unit sphere. Then, the average angular deviation of the normals is calculated within each bin. Finally, these per-bin statistics are assembled into the geometric descriptors.
@@
Following the spherical partitioning, the mean normal vector is computed for each sub-region within a cluster. The angular deviation between all unique pairs of these regional mean normal vectors within the same cluster is then calculated, generating $\binom{n}{2}$ distinct angular measurements per cluster (where $n$ is the number of sub-regions, e.g., $n=72$). A histogram of these angular deviations is constructed per cluster using 10-degree bin intervals (resulting in $d=18$ dimensions), yielding a $d$-dimensional geometric descriptor vector $\mathbf{F_{C}}\in\mathbb{R}^{d}$ for each cluster $C$.
@@
Subsequently, for each submap, the geometric descriptors $\mathbf{F_{C}}$ from all its clusters are collected and normalized. The VLAD algorithm [74] is then applied to this normalized set of descriptors to generate a single submap-level geometric descriptor. To accommodate submaps with varying cluster counts ($N_{c}$), we dynamically adjust the size of the VLAD codebook (i.e., the number of cluster centers, $K$) during encoding, setting $K$ proportionally to $N_{c}$. Submaps with a small $N_{c}$ use a correspondingly smaller $K$, while those with a large $N_{c}$ use a larger $K$. The resulting VLAD vector is then zero-padded to achieve a fixed dimensionality $a$. Finally, this yields the standardized geometric descriptor $\mathbf{F^{G}}\in\mathbb{R}^{a}$ for the submap. The entire process is illustrated in Fig. 3.
@@ III-D BEV descriptor
  1. 1)
    Log-Gabor filters: We project the acquired occupancy grid information onto the z-plane to obtain a BEV. The size of each pixel in the BEV corresponds to the physical size of the occupancy grids. Subsequently, we normalize the pixel values of the input image to the range [0, 255] to ensure numerical stability in subsequent processing. Then, we convert the input image’s coordinates (u,v) into polar coordinate representation $\left(\rho,\theta\right)$.Subsequently, we construct a Log-Gabor filter [75], defined as follows:
@@
  2. 2)
    Maximum Index Map (MIM): MIM is an orientation map that records the direction of the maximum response of each pixel in all directions [76]. That is, for each pixel $(\rho,\theta)$, find the direction $o$ with the maximum response:
@@ IV EXPERIMENTAL RESULTS AND ANALYSES
In this section, Section IV-A describes the dataset selected by us and some experimental settings, Section IV-B introduces the evaluation metrics of the experiment, Section IV-C presents the experimental results. Section IV-D and Section IV-E show the ablation experiment and running time situation respectively.
@@ IV-A Dataset and Experimental Settings
Figure 4: Datasets for evaluation.
@@
    MulRan Dataset [80]: This dataset contains data collected from multiple scenarios in South Korea using an Ouster OS1-64 LiDAR. We selected sequences 01 and 02 from the DCC, KAIST, and Riverside scenes for the experiment; these were labeled MulRan01 to MulRan06 respectively (e.g., DCC01=MulRan01, DCC02=MulRan02, etc.). Additionally, sequences Sejiong01 to Sejiong03 were used to conduct cross-temporal location recognition experiments.
Our feature extraction method is manually applied to the sub-maps derived from the accumulated point clouds. To benchmark our approach, we selected five handcrafted feature extraction methods for comparison on the evaluated datasets. These methods are: Scan Context++ [56], M2DP [54], NDT [81], BoW3D [82], and STD [18]. Among them, M2DP, NDT, BoW3D, and STD can be applied directly to the accumulated sub-maps. To ensure a fair comparison, we also applied the Scan Context++ method to the sub-maps. For this purpose, we re-projected the accumulated sub-map points onto the middle frame of the sub-map to simulate a dense ”scan” captured from that pose. For BoW3D, as its default configuration is tailored to the KITTI dataset, we only conducted experiments on KITTI and KITTI-360. Experiments involving our method and five handcrafted feature-based baseline approaches were executed on a laptop equipped with an Intel Core i7-10875H CPU @ 5.10GHz and 16GB RAM. For constructing the neural implicit representations, we used an RTX 2060 GPU for acceleration. The parameter values used in our method are listed in Table I.
TABLE I: Parameters of our methold
@@
$min\_size$ 200 minimum number of clusters
@@ IV-B Performance Evaluation Metrics
In this article, the conditions for establishing a loop closure are defined as follows: For a query sub-map, the average position of all its frames is computed and regarded as the position of that sub-map. If the distance between this sub-map’s position and that of a previous sub-map is less than or equal to 20 meters, and the difference in their indices is greater than 50, then these two sub-maps are considered positive (loop closure) pairs. The distance threshold of 20 meters was chosen based on considerations of the urban environment layout, typical LiDAR sensor range, and overall scale. Each method outputs a candidate frame along with its similarity score and compares the score against its predefined decision threshold. If the score exceeds the threshold, the match is classified as positive; otherwise, it is classified as negative. For a match classified as positive, the actual geometric distance between the query frame and the candidate frame is computed. If this distance is less than 20 meters, it is classified as a true positive (TP); otherwise, it is classified as a false positive (FP). For a match classified as negative: If no ground truth loop closure exists for the query frame (according to the criteria above), it is classified as a true negative (TN). However, if a ground truth loop closure does exist for the query frame, it is classified as a false negative (FN).
@@
    AUC: The Area Under the Curve (AUC) quantifies the area beneath the Precision-Recall (PR) curve. It serves as a widely adopted metric for evaluating model performance across various decision thresholds. As a single numerical value, AUC ranges from 0 to 1. A higher AUC score indicates superior performance, reflecting the model’s enhanced capacity to accurately classify positive instances across different threshold settings.
@@
    Max F1 Score: The F1 score, calculated using Eq 15, represents the harmonic mean of precision and recall. This score provides a balanced measure that combines both metrics. The Maximum F1 score serves as a comprehensive performance indicator, often used to identify the optimal decision threshold. By balancing the often-competing demands of precision and recall, it effectively reconciles these priorities. The peak F1 score value indicates the algorithm’s overall best performance, demonstrating an optimal equilibrium between correctly identifying true positives and minimizing false positives.
@@
Figure 5: Evaluation on twenty short-term sequences reveals that each subfigure illustrates the precision-recall performance of different methods. Our approach consistently outperforms all others across these sequences, demonstrating its robustness and adaptability.
TABLE II: AUC AND MAX F1 SCORES ON TESTING SEQUENCES
Method KITTI00 KITTI02 KITTI05 KITTI06 KITTI07 KITTI08 KITTI360-00 KITTI360-04 KITTI360-05 KITTI360-06
M2DP 0.64/0.65 0.31/0.45 0.55/0.47 0.52/0.58 0.28/0.43 0.43/0.58 0.48/0.56 0.08/0.23 0.09/0.28 0.18/0.31
NDT 0.69/0.70 0.52/0.59 0.61/0.67 0.6/0.68 0.38/0.52 0.52/0.6 0.51/0.57 0.26/0.44 0.44/0.56 0.39/0.48
SC++ 0.72/0.75 0.63/0.67 0.55/0.65 0.69/0.73 0.47/0.60 0.27/0.44 0.61/0.64 0.58/0.47 0.35/0.45 0.42/0.48
BoW3D 0.78/0.80 0.74/0.74 0.78/0.76 0.55/0.63 0.66/0.70 0.63/0.64 0.67/0.68 0.79/0.8 0.71/0.7 0.57/0.62
STD 0.51/0.58 0.85/0.8 0.78/0.76 0.71/0.77 0.57/0.62 0.73/0.74 0.71/0.71 0.6/0.64 0.68/0.71 0.63/0.64
Ours 0.93/0.86 0.95/0.88 0.93/0.91 0.91/0.89 0.91/0.82 0.83/0.78 0.89/0.86 0.87/0.85 0.87/0.86 0.76/0.77
Method NCLT01 NCLT02 NCLT03 NCLT04 MulRan01 MulRan02 MulRan03 MulRan04 MulRan05 MulRan06
M2DP 0.37/0.44 0.58/0.62 0.47/0.57 0.49/0.53 0.10/0.30 0.07/0.21 0.10/0.23 0.02/0.10 0.02/0.13 0.09/0.23
NDT 0.75/0.69 0.18/0.26 0.57/0.61 0.51/0.61 0.41/0.55 0.21/0.38 0.17/0.34 0.13/0.30 0.18/0.35 0.15/0.30
SC++ 0.74/0.74 0.67/0.64 0.63/0.69 0.44/0.52 0.41/0.55 0.36/0.48 0.34/0.48 0.24/0.42 0.29/0.44 0.17/0.32
STD 0.61/0.65 0.51/0.62 0.74/0.75 0.64/0.70 0.55/0.62 0.43/0.55 0.36/0.50 0.34/0.48 0.49/0.63 0.32/0.44
Ours 0.92/0.85 0.77/0.78 0.83/0.81 0.82/0.78 0.70/0.73 0.70/0.69 0.50/0.58 0.40/0.56 0.59/0.67 0.47/0.55
Note: On the left is the AUC and on the right is the max F1 score. The best-performing method is bolded.

### v3

@@ III-B Implicit Submap Representation With Elastic Neural Points
We will employ a memory-efficient implicit neural representation—a compact, learned model—to store the geometry of the sub-map. Based on elastic neural points (which means learned points that dynamically adjust), this representation possesses local rotational and translational invariance (remains the same when rotated or translated). Additionally, it is unaffected by the density of the original input point cloud (i.e., the number of points per space), has a uniform distribution (where points are spread out evenly), and can remove interference from certain dynamic objects (i.e., objects that move in the scene). Here, “elastic” signifies that the neural points undergo continual re-parameterisation: their 3-D positions and associated features are dynamically updated throughout sub-map accumulation (as the map grows, both their positions and features change).
1) Elastic Implicit Neural Point Cloud: The results of the sub-map composed of neural implicit points $\mathcal{M}$ are shown as follows:
$\mathcal{M}=\left\{\mathbf{m_{i}}=\left(\mathbf{p_{i}},\mathbf{q_{i}},\mathbf{f_{i}},t_{i}^{c},t_{i}^{u},\mu_{i}\right)\mid i=1,\dots,N\right\}$ (1)
@@
2) SDF Decoder: The joint feature coordinate encoding for scene representation is employed to achieve fast convergence and reconstruct surface holes. To ensure invariance of the prediction to the local translation and rotation of neural points, the query position is transformed from the global frame to the local frame of each neural point. Specifically, let $\mathbf{d_{j}}\in\mathbb{R}^{3}$ denote the query position in the local frame of neural point $\mathbf{m_{j}}$. This local coordinate $\mathbf{d_{j}}$ is obtained by transforming the global coordinates $\mathbf{p}_{s}$ of the query position using the pose of neural point $\mathbf{m_{j}}$, as follows:
@@
$D_{\theta}$ is a shallow MLP containing $M_{mlp}$ hidden layers with $N_{mlp}$ neurons.
Upon acquisition of a fresh LiDAR point cloud, each point is examined for eligibility as a new neural point. A point is instantiated as a neural point only if one of the following two mutually exclusive conditions is satisfied: 1) The corresponding entry in the hash table retains its default value, indicating the absence of any previously stored neural point; a new neural point is then initialized. 2) The hash-table entry is occupied, yet the Euclidean distance between the queried point p and the stored neural point exceeds a predefined threshold, signifying a hash collision. In this case, the incumbent neural point is evicted and replaced by a newly initialized one.
We employ a dynamic point filtering method based on the stability $\mu_{i}$ and SDF values of neural points. A point is classified as a dynamic point and filtered if its SDF value exceeds a dynamic distance threshold $\gamma_{d}$ and its stability value exceeds a stability threshold $\gamma_{\mu}$, Specifically, a sampled point $\mathbf{p_{W}}$ in the world coordinate system is considered dynamic if it satisfies conditions $\mathbf{S}(\mathbf{p_{W}})>\gamma_{d}$ and $\mathbf{H}(\mathbf{p_{W}})>\gamma_{\mu}$, where $\mathbf{S}(\mathbf{p_{W}})$ represents the SDF value indicating the distance from the point to the nearest surface and $\mathbf{H}(\mathbf{p_{W}})$ represents the stability value quantifying the consistency of the point across multiple observations. This method effectively removes dynamic objects from the map, enhancing its accuracy and consistency.
@@ III-C Normal Vector Descriptor
After getting the implicit submap representation, we extract the scene’s occupancy, the mesh’s points, and their normal directions.
To extract this information, the process begins by discretizing the submap into a voxel grid with a predefined resolution $r$. At each voxel corner point within this grid, we query the Signed Distance Function (SDF) value according to the method described previously. Based on these SDF values, we then apply an enhanced Marching Cubes (MC) algorithm to reconstruct a triangular mesh. This algorithm generates vertices uniformly distributed on the reconstructed surface. Subsequently, we calculate the normal vector for each vertex, defining the local surface orientation. To further refine the mesh, we filter the vertices to remove isolated outliers, thereby retaining only significant vertices along with their normals. Finally, by analysing the SDF values at the corners of each voxel, we derive occupancy information: voxels exhibiting a sign change in their corner SDF values are classified as occupied. This occupancy classification is essential for distinguishing between occupied and free space in the environment.
Subsequently, inspired by the work of FEC [72, 73], we apply a fast clustering method for mesh vertices. The algorithm is as Algorithm 1:
@@
1:mesh vertices point cloud $\mathbf{cloud}$, minimum cluster size min_sizemin\_size, distance tolerance toltol
@@
[…]
5:if n<min_sizen<min\_size then
@@
15:  if neighborsneighbors is empty then
@@
29:   if labels[j]>min_taglabels[j]>min\_tag then
30:     Relabel all points with $labels[k]=labels[j]$ to min_tagmin\_tag
@@
Following the derivation of clusters $\mathbf{\mathcal{C}=\left\{C_{1},...,C_{M}\right\}}$, points within each cluster are normalized. This normalization standardizes the scale of each cluster, thereby reducing sensitivity to noise and enhancing the robustness of subsequent geometric feature extraction. Specifically, all points in a cluster are isotropically scaled to be contained within a unit sphere of radius 1.
@@
Following normalization, the next step involves partitioning the spherical domain into 72 sub-regions based on longitude and latitude. The sphere is first divided along the z-axis into northern and southern hemispheres. By latitude, each hemisphere is further divided into two bands, resulting in 4 latitude bands globally. Longitudinal division then subdivides each latitude band into 18 sectors, formed by meridians spaced 20 degrees apart. Fig. 3 illustrates this spatial subdivision.
@@
Figure 3: (a) The figure illustrates the extraction of geometric descriptors for 3D segments. The input is a 3D segment with surface normals; color coding indicates normal deviations. First, these normals are mapped to discrete bins on a unit sphere, normalizing their distribution for systematic analysis. Then, the average angular deviation of the normals is calculated within each bin. Finally, these per-bin statistics are assembled into the geometric descriptors. (b) The figure illustrates the impact of keypoints on dominant 3-D segments, where diamond markers denote macro-level keypoints and colored spheres represent the distinct constituent points at the micro-level. The cardinality of keypoints assigned to each dominant segment subsequently governs the number of angular-distribution evaluations performed for that segment.
@@
Following the spherical partitioning, the mean normal vector is computed for each sub-region within a cluster. Next, the angular deviation between all unique pairs of these regional mean normal vectors within the same cluster is calculated. This process generates $\binom{n}{2}$ distinct angular measurements per cluster, where $n$ is the number of sub-regions (e.g., $n=72$). We then randomly select $n_{m}$ measurements. A histogram of these angular deviations is constructed per cluster using 10-degree bin intervals, resulting in $d=18$ dimensions. This process yields a $d$-dimensional geometric descriptor vector $\mathbf{F_{C}}\in\mathbb{R}^{d}$ for each cluster $C$.
@@
Subsequently, for each 3D segment, the geometric descriptors $\mathbf{F_{C}}$ from all its clusters are collected and normalised. The VLAD algorithm [74] is then applied to this normalised set of descriptors to generate a single submap-level geometric descriptor. To accommodate submaps with varying cluster counts ($N_{c}$), we dynamically adjust the size of the VLAD codebook (i.e., the number of cluster centres, $K$) during encoding, setting $K$ proportionally to $N_{c}$. Submaps with a small $N_{c}$ use a correspondingly smaller $K$, while those with a large $N_{c}$ use a larger $K$. The resulting VLAD vector is then zero-padded to achieve a fixed dimensionality $a$. Finally, this yields the standardised geometric descriptor $\mathbf{F^{G}}\in\mathbb{R}^{a}$ for the submap. The entire process is illustrated in Fig. 3.
For each primary 3-D segment, we dynamically schedule the computation of micro-geometric descriptors. The number of evaluations depends on how many BEV feature points project into the segment. Specifically, the descriptor is computed between 1 and 10 times, using a linear mapping from the point count. Fig. 3 shows a diagram of this procedure. Details on BEV feature extraction are in Section III-D. Additionally, we adjust the computation frequency of the angular distribution for a segment based on the number of macro-level keypoints assigned to it. This design is motivated by observing that regions with many feature points often encode salient geometric structures. Allocating more computation to these regions yields more distinctive descriptors.
@@ III-D BEV descriptor
1) Log-Gabor filters: We project the acquired occupancy grid information onto the z-plane to obtain a BEV. During projection, we retain only points found near each dominant 3-D segment $\mathbf{\mathcal{C}=\left\{C_{1},...,C_{M}\right\}}$. This effectively suppresses spurious and noisy observations. At an early and efficient clustering stage, road-surface points are discarded. We prune these because the implicit surface representation generates angular vertices that are sparsely scattered in geometrically impoverished regions. As a result, they fall below the saliency threshold and are removed during the clustering process. The size of each pixel in the BEV matches the physical size of the occupancy grids. Next, we normalise the pixel values of the input image to the range [0, 255] to ensure numerical stability in subsequent processing. We then convert the input image’s coordinates (u, v) into polar coordinates $(\rho,\theta)$. Finally, we construct a Log-Gabor filter [75], defined as follows:
@@
The preset scale $1,\dots,N_{s}. ss$ represents the scale index of the Log-Gabor filter.
2) Maximum Index Map (MIM): MIM is an orientation map that records the direction of the maximum response of each pixel in all directions [76]. That is, for each pixel $(\rho,\theta)$, find the direction $o$ with the maximum response:
@@ IV EXPERIMENTAL RESULTS AND ANALYSES
In this section, Section IV-A describes the dataset selected by us and some experimental settings, Section IV-B introduces the evaluation metrics of the experiment, and Section IV-C presents the experimental results. Section IV-D and Section IV-E show the ablation experiment and running time and storage efficiency, respectively.
@@ IV-A Dataset and Experimental Settings
Figure 4: Datasets for experiment.
@@
    MulRan Dataset [80]: This dataset contains data collected from multiple scenarios in South Korea using an Ouster OS1-64 LiDAR. We selected sequences 01 and 02 from the DCC, KAIST, and Riverside scenes for the experiment; these were labeled MulRan01 to MulRan06, respectively (e.g., DCC01=MulRan01, DCC02=MulRan02, etc.). Additionally, sequences Sejiong01 to Sejiong03 were used to conduct cross-temporal location recognition experiments.
Our feature extraction method is manually applied to the sub-maps derived from the accumulated point clouds. To benchmark our approach, we selected five handcrafted feature extraction methods for comparison. These were evaluated on the selected datasets. The methods are Scan Context++ [56], M2DP [54], NDT [81], BoW3D [82], Ring++ [83], and STD [18]. M2DP, NDT, BoW3D, and STD can be applied directly to the accumulated sub-maps. To ensure a fair comparison, we also applied the Scan Context++ and Ring++ methods to the sub-maps. For this purpose, we re-projected the accumulated sub-map points onto the middle frame of the sub-map. This step simulated a dense ”scan” captured from that pose. BoW3D’s default configuration targets the KITTI dataset; therefore, we conducted our experiments only on the KITTI and KITTI-360 datasets. Ring++’s default configuration is tailored to the NCLT and Mulran datasets; therefore, we used those datasets exclusively for Ring++. All experiments were conducted on a laptop equipped with an Intel Core i7-10875H CPU at 5.10 GHz and 16GB RAM. We utilized an RTX 2060 GPU for acceleration during the construction of the neural implicit representations. The parameter values used in our method are listed in Table I.
TABLE I: Parameters of our method
@@
$\gamma_{\mu}$ 4 Dynamic filtering threshold for stability
$M_{nlp}$, $N_{nlp}$ 8 MLP level and neuron count
@@
min_sizemin\_size 200 Minimum number of clusters
@@
$n_{m}$ 400 Number of angle differences
$N_{s}$ 4 Number of scales of Log-Gabor filter
@@ IV-B Performance Evaluation Metrics
In this article, the conditions for establishing a loop closure are defined as follows. For a query sub-map, first compute the average position of all its frames to represent the sub-map’s position. If the distance between this position and that of a previous sub-map is less than or equal to 20 meters, and their indices differ by more than 50, then these two sub-maps are considered positive (loop closure) pairs. The 20-meter distance threshold was chosen based on the urban environment, the typical range of LiDAR sensors, and the overall scale. Each method outputs a candidate frame with its similarity score and compares the score to a predefined decision threshold. If the score exceeds the threshold, classify the match as positive; if not, classify it as negative. For a match classified as positive, compute the actual geometric distance between the query frame and the candidate frame. If this distance is less than 20 meters, it is considered a true positive (TP); otherwise, it is classified as a false positive (FP). For a match classified as negative, if no ground truth loop closure exists for the query frame (according to the above criteria), it is considered a true negative (TN). If a ground truth loop closure exists for the query frame, it is a false negative (FN).
@@
    AUC: The Area Under the Curve (AUC) measures the area under the Precision-Recall (PR) curve. It is a common metric for checking model performance at different thresholds. AUC is a single number between 0 and 1. A higher AUC indicates that the model is more effective at identifying positive cases across various thresholds.
@@
    Max F1 Score: The F1 score, calculated using Eq 15, represents the harmonic mean of precision and recall. This score provides a balanced measure by combining both metrics. The Maximum F1 score acts as a comprehensive performance indicator. It is often used to identify the optimal decision threshold. By balancing the competing demands of precision and recall, it effectively reconciles these priorities. The peak F1 score value reflects the algorithm’s optimal performance. It demonstrates a balance between accurately identifying true positives and minimizing false positives.
@@
Figure 5: Evaluation of twenty short-term sequences reveals that each subfigure in Figure 5 corresponds to the precision-recall performance of a specific method on these sequences. Our approach consistently outperforms all others across these sequences, demonstrating its robustness and adaptability. All image coordinates (x, y) are normalized to the unit finterval [0, 1], and the subsequently figure PR curves adhere to the same convention.
TABLE II: Comprehensive Performance Evaluation Across Multiple Datasets (AUC $\uparrow$ / Max F1 Score $\uparrow$ )
Method KITTI KITTI-360
00
02
05
06
00
04
05
06
M2DP 0.64/0.65 0.31/0.45 0.55/0.47 0.52/0.58 0.48/0.56 0.08/0.23 0.09/0.28 0.18/0.31
NDT 0.69/0.70 0.52/0.59 0.61/0.67 0.60/0.68 0.51/0.57 0.26/0.44 0.44/0.56 0.39/0.48
SC++ 0.72/0.75 0.63/0.67 0.55/0.65 0.69/0.73 0.61/0.64 0.58/0.47 0.35/0.45 0.42/0.48
BoW3D 0.78/0.80 0.74/0.74 0.78/0.76 0.55/0.63 0.67/0.68 0.79/0.80 0.71/0.70 0.57/0.62
STD 0.51/0.58 0.85/0.80 0.78/0.76 0.71/0.77 0.71/0.71 0.60/0.64 0.68/0.71 0.63/0.64
Ours 0.93/0.86 0.95/0.88 0.93/0.91 0.91/0.89 0.89/0.86 0.87/0.85 0.87/0.86 0.76/0.77
Method NCLT MulRan
01
02
03
04
01
02
03
04
M2DP 0.37/0.44 0.58/0.62 0.47/0.57 0.49/0.53 0.10/0.30 0.07/0.21 0.10/0.23 0.02/0.10
NDT 0.75/0.69 0.18/0.26 0.57/0.61 0.51/0.61 0.41/0.55 0.21/0.38 0.17/0.34 0.13/0.30
SC++ 0.74/0.74 0.67/0.64 0.63/0.69 0.44/0.52 0.41/0.55 0.36/0.48 0.34/0.48 0.24/0.42
STD 0.61/0.65 0.51/0.62 0.74/0.75 0.64/0.70 0.55/0.62 0.43/0.55 0.36/0.50 0.34/0.48
Ring++ 0.78/0.77 0.72/0.71 0.77/0.79 0.77/0.79 0.63/0.66 0.57/0.59 0.42/0.53 0.35/0.53
Ours 0.92/0.85 0.77/0.78 0.83/0.81 0.82/0.78 0.70/0.73 0.70/0.69 0.50/0.58 0.40/0.56
Note: This table presents the Area Under the Curve (AUC) and maximum F1 scores for all evaluated methods across multiple datasets. The best result for each dataset is highlighted in bold.